Find the following probabilities based on the standard normal variable Z (Round your answers to 4 decimal places.) a. P(Z> 1.04) b. P(Zs -1.74) c. P(O s Z s 1.81) d. P(-0.81 s Zs 2.66)
Answer #1

Solution :

Using standard normal table,

(a)

P(Z > 1.04) = 1 - P(Z < 1.04) = 1 - 0.8508 = 0.1492

(b)

P(Z \leq -1.74) = 0.0409

(c)

P(0 \leq Z \leq 1.81) = P(Z \leq 1.81) - P(Z \leq 0) = 0.9649 - 0.5 = 0.4649

(d)

P(-0.81 \leq Z \leq 2.66) = P(Z \leq 2.66) - P(Z \leq -0.81) = 0.9961 - 0.2090 = 0.7871